OpenStudy (kanwal32):

the sum upto infinity of 1/1+1/1+2+1/1+2+3......

OpenStudy (kanwal32):

\[\frac{1}{1}+\frac{1}{1+2}+\frac{1}{1+2+3}......\]

OpenStudy (kanwal32):

@UnkleRhaukus help

OpenStudy (kanwal32):

@ganeshie8 hlp

OpenStudy (ikram002p):

ok so next term is 1/1+2+3+4 right ?

OpenStudy (kanwal32):

can we write last terms as zero

OpenStudy (ikram002p):

Hint :- \(1+2+3+... + n = \dfrac{n(n+1)}{2}\)

OpenStudy (ikram002p):

so the series we have would be like \(\sum_{i=1}^{n} \dfrac {1}{\dfrac{i(i+1)}{2}}\)

OpenStudy (kanwal32):

so1/ n(n+1) will be zero

OpenStudy (ikram002p):

mm limit when n goes to infinity =0

OpenStudy (kanwal32):

answer is 2

OpenStudy (kanwal32):

ok i got it thanks @ikram002p

OpenStudy (unklerhaukus):

triangle numbers

OpenStudy (ikram002p):

it will be inverse of triangle number like this :- \(2(\dfrac{1}{2} +\dfrac{1}{6} +\dfrac{1}{12} +\dfrac{1}{20} +...)\)

OpenStudy (ikram002p):

@kanwal32 np :)